OpenStudy (anonymous):

solve this systems of linear equations using any algebraic method. 3y=11-2x 3x=y-11

OpenStudy (lgbasallote):

3x = y - 11 therefore y = 3x + 11 substitute this into the first equation 3(3x + 11) = 11 - 2x 9x + 33 = 11 - 2x 9x + 2x = 11 - 33 11x = -22 x = -2

OpenStudy (lgbasallote):

if i substitute that to the second equation 3(-2) = y - 11 -6 = y - 11 -6 + 11 = y y = 5